# Privately Estimating Black-Box Statistics

### v1

@@ 1 Introduction
In many cases, the global sensitivity of the function we want to evaluate is large (or even infinite), or simply unknown. In practice, the function may be given to us as a “black box” – that is, we can only evaluate the function as a “oracle” and not inspect its inner workings, or it may be presented as a piece of untrusted code that is too complicated to analyse. In these cases, we cannot rely on the standard noise addition approach to ensure differential privacy. Many methods have been proposed for privately evaluating functions that go beyond noise addition and are applicable to functions with high (global) sensitivity; see Section 2 for a brief survey. However, all of these methods have various drawbacks (which we briefly discuss next) that have limited practical adoption: Either they require evaluating the function many times (or they require some a priori structural knowledge about the function), or they are statistically inefficient.
Most methods for evaluating functions with high global sensitivity still rely on being able to compute or bound some property of the function, such as smooth sensitivity [NRS07] or distance to instability [DL09], instead of global sensitivity. Hence these methods require either careful analysis of the function or evaluating of the function over a large fraction of its domain (i.e., they may require evaluating the function on exponentially many – or even infinitely many – inputs). This makes these methods impractical in the black-box setting.
@@
There is one method that does not have the aforementioned limitations: The sample-and-aggregate framework of [NRS07] only requires evaluating the function on a small number of inputs (with each input consisting only of “real” data) and does not require any structural assumptions about the function, which makes it appealing in practice [PAEGT17, PSMRTE18]. The catch is that this method is statistically inefficient in terms of its use of data. That is, if we start with a dataset of size $n$, then sample-and-aggregate evaluates the function on datasets of size $O(\varepsilon n)$. (Specifically, the sample-and-aggregate framework partitions the dataset into $O(1/\varepsilon)$ equal-sized parts.) Very roughly, the final accuracy of our private estimate given a dataset of size $n$ is only as good as a non-private estimate on a dataset of size $O(\varepsilon n)$. Thus, if the privacy parameter $\varepsilon$ is small, then sample-and-aggregate suffers a significant cost in terms of statistical accuracy.
@@ 1.1 Our Contributions
In this article, we examine the tradeoff between statistical efficiency (i.e., how much data is needed to estimate a statistic) and oracle query complexity (i.e., how many times do we need to evaluate the function).
Our main result is a differentially private algorithm which takes a real-valued black-box function $f$ and a private dataset $x$ and evaluates the function on multiple subsets of the input dataset and then outputs an estimate $y$ for the value of the function. Our algorithm interpolates between sample-and-aggregate [NRS07] and more recent computationally inefficient approaches [CD20, FDY22, LRSS25, etc.]
@@
Let [^1] We precisely define $C(n,m,t)$ later in Definition 3.6; for now the given upper bound suffices.
$k=C(n,m,t)<\frac{{n\choose t}}{{m\choose t}}\left(1+\log{m\choose t}\right)+1\leq O\left(\left(\frac{n-t+1}{m-t+1}\right)^{t}\cdot m\right).$ (1.2)
@@
  • Statistical Accuracy: Let $\mathcal{D}$ be an arbitrary probability distribution on $\mathcal{X}$. Suppose ${\underset{X\leftarrow\mathcal{D}^{n-m}}{\mathbb{P}}\left[|f(X)-\nu|\leq\alpha\right]}\geq 1-\beta$. Then ${\underset{X\leftarrow\mathcal{D}^{n}}{\mathbb{P}}\left[|M^{f}(X)-\nu|\leq\alpha\right]}\geq 1-k\beta$.
@@
      Setting $m=\frac{t}{t+1}n$ yields $n-m=\frac{n}{t+1}$ and k=t+1k=t+1, [^2] This value of $k=C(n,\frac{t}{t+1}n,t)\leq t+1$ is tighter than the upper bound given in the theorem statement; see Equation 3.31. which corresponds to sample-and-aggregate [NRS07]. That is, the cost of privacy is a multiplicative factor of $t$ in the sample complexity – i.e., with $n$ private samples we get accuracy comparable to $n-m=\Theta(n/t)$ non-private samples. This is the most computationally efficient instantiation of our result.
@@
      The above points (a and b) are the extremes on the tradeoff curve; now we consider a setting that interpolates between these: Setting $m=\frac{tn}{t+c}$ yields $k=O(t^{c})$. (Here $c\geq 1$ is an appropriate integer.) Relative to (a) sample-and-aggregate, this increases the number of data points $n-m=\frac{cn}{t+c}$ available for each evaluation by a factor of almost $c$ – i.e., we go from $n-m=\Theta(n/t)$ to $n-m=\Theta(cn/t)$ – at the expense of a polynomial increase in the number of evaluations $k$. This parameter setting is likely of practical interest.
@@
assuming $\delta\leq(\varepsilon/10)^{1.1}$.
@@ 1.2 Our Techniques
Our algorithm can be viewed as an extension of the sample-and-aggregate paradigm [NRS07]. Namely, we evaluate the function on subsets of the input and then we aggregate those values in a way that ensures differential privacy. The novelty is in how we choose the subsets and how we aggregate the values.
Our algorithm has two main technical ingredients: First we use a combinatorial object known as a covering design or a Turán system to select $k$ overlapping subsets of the input on which to evaluate the function. Second, we use a variant of the shifted inverse mechanism of [FDY22] to aggregate the values in a differentially private manner.
@@
It only remains to aggregate the values in a way that this form of robustness translates into differential privacy. Computing a differentially private mean or median of the $k$ values does not suffice, since a single datapoint could affect a majority of the values. This is where the shifted inverse mechanism fits in.
@@
To illustrate how the shifted inverse mechanism works, consider the special case with a binary output space $\mathcal{Y}=\{0,1\}$. (The general case can, with some loss, be reduced to this case.) Now we ask “how many input datapoints would I need to remove so that all of the remaining output values are all 0?” (That is, if we remove a datapoint, then all of the output values that depend on that datapoint are removed.) If all of the output values are 0, then the answer to the query is 0 – i.e., no datapoints need to be removed. If all of the output values are $1$, then the answer to the query is at least $t$, by the properties of the covering design. By construction, this query has sensitivity $1$; thus the answer to the query can be approximated in a differentially private manner by adding Laplace or Gaussian noise. As long as $t$ is large enough, we can accurately distinguish between the case where all of the output values are 0 and the case where all of them are $1$. (When some values are 0 and some are $1$, the outcome is indeterminate.) Assuming each value is individually accurate with high probability, the aggregated value will also be accurate with high probability.
@@
Our lower bound mirrors the intuition for our algorithm. That is, $(\varepsilon,\delta)$-differential privacy requires robustness to $t=O(\log(1/\delta)/\varepsilon)$ corruptions. To be precise, we perform a packing argument [HT10]. That is, we use group privacy to argue that changing $t$ values cannot totally change the output of the algorithm. We restrict the algorithm to evaluating the function on subsets of the input of size nmn-m. (We can force this restriction using standard tricks, such as making the function fail when given an input that isn’t of this form.) Now the only way for the algorithm to satisfy the given level of robustness is for it to query enough sets so that at least one of them contains none of the $t$ corrupted points. Roughly, this implies that the subsets queried by the algorithm must form a covering design. The lower bound then follows.
@@ 2.1 Alternatives to Global Sensitivity
The inverse sensitivity mechanism [AD20, MMNW11, JS13, Ste23] provides a loss function with low global sensitivity for any function of interest; the exponential mechanism [MT07] can then be applied to estimate the value of interest. The idea is simple: Given a function $f$, a dataset $x$, and a value $y$, define the loss $\ell(x,y)=\min\{|x\setminus x^{\prime}|+|x^{\prime}\setminus x|:f(x^{\prime})=y\}$ to be the least number of elements of $x$ that need to be added or removed in order the change the function value to $y$. Clearly, $\ell(x,y)=0$ if and only if $y=f(x)$. Furthermore, if the local sensitivity of $f$ at $x$ is small, then $\ell(x,y)\leq 1$ implies $y\approx f(x)$. More generally, if $f$ is appropriately well-behaved near $x$, then any approximate minimiser $y$ of the loss $\ell(x,y)$ must be a good approximation to $f(x)$.
Another approach is to replace the function of interest with a function that has low global sensitivity and which still provides a good approximation to the original function. For example, when computing a sum, we might clip the values to ensure that they are bounded; if the clipping threshold is chosen appropriately, this ensures low global sensitivity and doesn’t change the value of the function too much. In general, Lipschitz extensions provide low-sensitivity approximating functions that can be used in the differentially private setting [KNRS13, BBDS13, RS15, RS16a, RS16].
@@
All of the aforementioned methods suffer from computational intractability. They are only practical in special cases where we can analytically compute the relevant bounds. The underlying reason for this is that the quantities of interest all involve some universal quantification over datasets. In general, implementing these methods requires enumerating exponentially many inputs – or even infinitely many. In particular, they are not practical for functions that are given to us as a black box.
@@ 2.2 Down-Local Algorithms
Recent work [CD20, FDY22, KL23, LRSS25] has sought to overcome the aforementioned limitations by devising algorithms that are down-local in the sense that, given a function $f$ and a dataset $x$, the algorithm only evaluates $f$ on subsets of $x$, rather than on arbitrary points in its domain.
In particular, down-local algorithms only evaluate the function on “real” data; this is an added benefit, since many functions might “break” when given arbitrary inputs. For example, even the mean has infinite local sensitivity, when we allow unbounded inputs, even though it may be well-behaved for real inputs.
@@
Sample-and-aggregate [NRS07] is closely related to our approach. In its simplest form, [^3] More generally, sample-and-aggregate allows overlapping subdatasets, but then the aggregator must handle the fact that a single person’s data may affect multiple function values. sample-and-aggregate partitions the dataset into smaller subdatasets, evaluates the function of interest on each subdataset, and then aggregates the function values in a differetially private manner (e.g., using smooth sensitivity applied to the aggregation function). A single person’s data will be in only one of the subdatasets and so a single person can only affect one of the values. The advantage of the sample and aggregate approach is that requires no structural knowledge about the function of interest and is computationally efficient. (Intuitively, it pushes the privacy analysis onto the aggregation function, rather than the function we want to evaluate.) The downside is that we evaluate the function on the smaller subdatasets. This can lead to significant loss in accuracy relative to (non-privately) evaluating the function on the whole dataset. Our algorithm addresses the downside of sample and aggregate by allowing us to evaluate the function on larger subdatasets, at the expense of requiring us to evaluate on more of these subdatasets.

### v2

@@ 1 Introduction
In many cases, the global sensitivity of the function we want to evaluate is large (or even infinite), or simply unknown. In practice, the function may be given to us as a “black box” – that is, we can only evaluate the function as a “oracle” and not inspect its inner workings, or it may be presented as a piece of untrusted code that is too complicated to analyse. [^1] As a motivating practical example of code that is too complicated to analyze, consider $f$ such that computing $y=f(x)$ requires training a machine learning model with the input $x$ being the training data and the output $y$ being the predicted label of a fixed unlabelled data point [PAEGT17, PSMRTE18]. In these cases, we cannot rely on the standard noise addition approach to ensure differential privacy. Many methods have been proposed for privately evaluating functions that go beyond noise addition and are applicable to functions with high (global) sensitivity; see Section 2 for a brief survey. However, all of these methods have various drawbacks (which we briefly discuss next) that have limited practical adoption: Either they require evaluating the function many times (or they require some a priori structural knowledge about the function), or they are statistically inefficient.
Most methods for evaluating functions with high global sensitivity still rely on being able to compute or bound some property of the function, such as smooth sensitivity [NRS07] or distance to instability [DL09], instead of global sensitivity. Hence these methods require either careful analysis of the function or evaluating the function over a large fraction of its domain (i.e., they may require evaluating the function on exponentially many – or even infinitely many – inputs). This makes these methods impractical in the black-box setting.
@@
There is one method that does not have the aforementioned limitations: The sample-and-aggregate framework of [NRS07] only requires evaluating the function on a small number of inputs (with each input consisting only of “real” data) and does not require any structural assumptions about the function, which makes it appealing in practice [PAEGT17, PSMRTE18, CCWLNSS23]. The catch is that this method is statistically inefficient in terms of its use of data. That is, if we start with a dataset of size $n$, then sample-and-aggregate evaluates the function on datasets of size $O(\varepsilon n)$. (Specifically, the sample-and-aggregate framework partitions the dataset into $O(1/\varepsilon)$ equal-sized parts.) Very roughly, the final accuracy of our private estimate given a dataset of size $n$ is only as good as a non-private estimate on a dataset of size $O(\varepsilon n)$. Thus, if the privacy parameter $\varepsilon$ is small, then sample-and-aggregate suffers a significant cost in terms of statistical accuracy.
@@ 1.1 Our Contributions
In this article, we examine the tradeoff between statistical efficiency (i.e., how much data is needed to estimate a statistic) and oracle query complexity (i.e., how many times we need to evaluate the function).
Our main result is a differentially private algorithm which takes a real-valued black-box function $f$ and a private dataset $x$ and evaluates the function on multiple subsets of the input dataset and then outputs an estimate $y$ for the value of the function. Our algorithm interpolates between sample-and-aggregate [NRS07] and the more recent computationally inefficient algorithm of [LRSS25].
@@
Let [^2] We precisely define $C(n,m,t)$ later in Definition 3.6; for now the given upper bound suffices. Throughout, $\log$ denotes the natural logarithm with base $e\approx 2.718$.
$k=C(n,m,t)<\frac{{n\choose t}}{{m\choose t}}\left(1+\log{m\choose t}\right)+1\leq O\left(\left(\frac{ne}{m}\right)^{t}\cdot t\log m\right).$ (1.2)
@@
  • Statistical Accuracy: Let $\mathcal{D}$ be an arbitrary probability distribution on $\mathcal{X}$. Suppose ${\underset{X\leftarrow\mathcal{D}^{n-m}}{\mathbb{P}}\left[|f(X)-\nu|\leq\alpha\right]}\geq 1-\beta$ for some $\alpha,\beta,\nu\in\mathbb{R}$. Then ${\underset{X\leftarrow\mathcal{D}^{n}}{\mathbb{P}}\left[|M^{f}(X)-\nu|\leq\alpha\right]}\geq 1-k\beta$.
@@
      Setting $m=\frac{t}{t+1}n$ yields $n-m=\frac{n}{t+1}$ and $k\leq t+1$, [^3] This value of $k=C(n,\frac{t}{t+1}n,t)\leq t+1$ is tighter than the upper bound given in the theorem statement; see Equation 3.16. For this informal discussion, we ignore integer divisibility issues. which corresponds to sample-and-aggregate [NRS07]. That is, the cost of privacy is a multiplicative factor of $t$ in the sample complexity – i.e., with $n$ private samples we get accuracy comparable to $n-m=\Theta(n/t)$ non-private samples. This is the most computationally efficient instantiation of our result.
@@
      The above points (a and b) are the extremes on the tradeoff curve; now we consider a setting that interpolates between these: Setting $m=\frac{tn}{t+c}$ yields $k\leq O(\min\{e^{t+c}\cdot t\log m,t^{c}\})$. (Here $c\geq 1$ is an appropriate integer.) Relative to (a) sample-and-aggregate, this increases the number of data points $n-m=\frac{cn}{t+c}$ available for each evaluation by a factor of almost $c$ – i.e., we go from $n-m=\Theta(n/t)$ to $n-m=\Theta(cn/t)$ – at the expense of a modest increase in the number of evaluations $k$. This parameter setting is likely of practical interest.
@@
assuming $\delta\leq(\varepsilon/10)^{1.1}\leq 1$.
@@ 1.2 Our Techniques
Algorithm: Our algorithm can be viewed as an extension of the sample-and-aggregate paradigm [NRS07]. Namely, we evaluate the function on subsets of the input and then we aggregate those values in a way that ensures differential privacy. The novelty is in how we choose the subsets and how we aggregate the values.
Our algorithm has two main technical ingredients: First we use a combinatorial object known as a covering design or a Turán system to select $k$ overlapping subsets of the input on which to evaluate the function. Second, following [LRSS25], we use a variant of the shifted inverse mechanism of [FDY22] to aggregate the values in a differentially private manner.
@@
It only remains to aggregate the values in a way that translates this form of robustness into differential privacy. Computing a differentially private mean or median of the $k$ values does not suffice, since a single datapoint could affect a majority of the values. This is where the shifted inverse mechanism fits in.
@@
To illustrate how the shifted inverse mechanism works, consider the special case with a binary output space $\mathcal{Y}=\{0,1\}$. (The general case can, with some loss, be reduced to this case.) Now we ask “how many input datapoints would I need to remove so that the remaining output values are all 0?” (That is, if we remove a datapoint, then all of the output values that depend on that datapoint are removed.) If all of the output values are 0, then the answer to the query is 0 – i.e., no datapoints need to be removed. If all of the output values are $1$, then the answer to the query is at least $t$, by the properties of the covering design. By construction, this query has sensitivity $1$; thus the answer to the query can be approximated in a differentially private manner by adding Laplace or Gaussian noise. As long as $t$ is large enough, we can accurately distinguish between the case where all of the output values are 0 and the case where all of them are $1$. (When some values are 0 and some are $1$, the outcome is indeterminate.) Assuming each value is individually accurate with high probability, the aggregated value will also be accurate with high probability.
Example Instantiation: To give further intuition for how our algorithm works, we consider instantiating it with a particular simple covering design. Again consider a binary output space $\mathcal{Y}=\{0,1\}$ and let $t=O\big(\frac{1}{\varepsilon}\log(1/\delta)\big)$. Partition the dataset (of size $n$) into $t+2$ parts of (almost) equal size; call these $D_{1},\cdots,D_{t+2}$. Now we evaluate the function $f:\mathcal{X}^{*}\to\mathcal{Y}$ on all $\frac{(t+1)(t+2)}{2}$ pairs $D_{i}\cup D_{j}$. That is, we evaluate the function on subdatasets of size $2\frac{n}{t+2}$; so each pair gets a label of 0 or $1$. (Note that partitioning into $t+2$ equal-sized parts and considering the union of pairs of parts corresponds to a $(n,m,t)$-covering design (Definition 3.6) with $m=2\frac{n}{t+2}$.)
Now we apply the shifted inverse mechanism, which entails computing the minimum number of parts $D_{1},\cdots,D_{t+2}$ that need to be chosen to cover all pairs that have a label of $1$. If all of the pairs have a label of 0, then this number is zero; if all of the pairs have a label of $1$, then this number is $t+1$. Since this number is low sensitivity, it can be privately approximated, which yields the final output.
Note that computing this number corresponds to computing the size of a minimum vertex cover of the following graph: Each part $D_{i}$ corresponds to a vertex and, for each pair $i,j$, there is an edge between $D_{i}$ and $D_{j}$ if and only $f(D_{i}\cup D_{j})=1$. Since vertex cover is NP-complete [Kar72], computing this value is potentially intractable. See Section 6.2 for further discussion of this limitation.
Comparing this example instantiation with sample-and-aggregate, we see that we evaluate the function $f$ on datasets of size approximately $2n/t$ rather than $n/t$, which yields better statistical accuracy. In exchange, we need approximately $t^{2}/2$ evaluations of the function $f$, rather than $t$. This is one instantiation the tradeoff that our algorithm offers.
@@
Lower Bound: Our lower bound mirrors the intuition for our algorithm. That is, $(\varepsilon,\delta)$-differential privacy requires robustness to $t=O(\log(1/\delta)/\varepsilon)$ corruptions. To be precise, we perform a packing argument [HT10]. That is, we use group privacy to argue that changing $t$ values cannot totally change the output of the algorithm. We restrict the algorithm to evaluating the function on subsets of the input of size nmn-m. (We can force this restriction using standard tricks, such as making the function fail when given an input that isn’t of this form.) Now the only way for the algorithm to satisfy the given level of robustness is for it to query enough sets so that at least one of them contains none of the $t$ corrupted points. Roughly, this implies that the subsets queried by the algorithm must form a covering design. The lower bound then follows.
@@ 2 Related Work
There is a long line of work on differentially privately evaluating a function on a sensitive dataset that makes weaker and weaker assumptions on the underlying function, which we briefly survey in this section. [JR13] were the first to explicitly study the black-box setting, where we make no assumptions about the function beyond what we can glean from evaluating it.
@@ 2.1 Alternatives to Global Sensitivity
The inverse sensitivity mechanism [AD20, MMNW11, JS13, Ste23, AUZ23, HKMN23] provides a loss function with low global sensitivity for any function of interest; the exponential mechanism [MT07] can then be applied to estimate the value of interest. The idea is simple: Given a function $f$, a dataset $x$, and a value $y$, define the loss $\ell(x,y)=\min\{|x\setminus x^{\prime}|+|x^{\prime}\setminus x|:f(x^{\prime})=y\}$ to be the least number of elements of $x$ that need to be added or removed in order the change the function value to $y$. Clearly, $\ell(x,y)=0$ if and only if $y=f(x)$. Furthermore, if the local sensitivity of $f$ at $x$ is small, then $\ell(x,y)\leq 1$ implies $y\approx f(x)$. More generally, if $f$ is appropriately well-behaved near $x$, then any approximate minimiser $y$ of the loss $\ell(x,y)$ must be a good approximation to $f(x)$.
Another approach is to replace the function of interest with a function that has low global sensitivity and which still provides a good approximation to the original function. For example, when computing a sum, we might clip the values to ensure that they are bounded; if the clipping threshold is chosen appropriately, this ensures low global sensitivity and doesn’t change the value of the function too much. In general, Lipschitz extensions provide low-sensitivity approximating functions that can be used in the differentially private setting [KNRS13, BBDS13, RS15, RS16a, RS16]. In particular, a local Lipschitz filter [JR13, AJMR15, LLRV25] provides a black-box method for constructing a Lipschitz extension.
@@
All of the aforementioned methods (except sample-and-aggregate) suffer from computational intractability. They are only practical in special cases where we can analytically compute the relevant bounds. The underlying reason for this is that the quantities of interest all involve some universal quantification over datasets. In particular, they are not practical for functions that are given to us as a black box, as they would require enumerating exponentially many inputs – or even infinitely many.
An additional limitation is that all of the above methods (except sample-and-aggregate) fail when the function $f$ has high local sensitivity at the given dataset $x$. That is, there exists some input $x^{\prime}$ neighbouring the given input $x$ such that the function value on that input $f(x^{\prime})$ is far from the actual function value $f(x)$. In particular, adding one corrupted data point to the given input $x$ might “break” the function – i.e. change the function value arbitrarily. For example, even the mean has infinite local sensitivity, when we allow unbounded inputs, even though it may be well-behaved for realistic inputs. (Note that high local sensitivity is an issue even if we are not in the black-box setting.)
@@ 2.2 Down-Local Algorithms
Recent work [CZ13, RS16, CD20, FDY22, KL23, LRSS25] has sought to overcome the aforementioned limitations by devising algorithms that are down-local in the sense that, given a function $f$ and a dataset $x$, the algorithm only evaluates $f$ on subsets of the given dataset $x$, rather than on arbitrary points in its domain.
This has two benefits: First, it restricts the number of inputs on which we can evaluate a black-box function $f$. Second, down-local algorithms only evaluate the function on “real” data, which avoids the problem that some functions might “break” when given arbitrary inputs. In other words, down-local algorithms work well with functions that have high local sensitivity, but are well-behaved on realistic inputs.
@@
Sample-and-aggregate [NRS07] is closely related to our approach. In its simplest form, [^4] More generally, sample-and-aggregate allows overlapping subdatasets, but then the aggregator must handle the fact that a single person’s data may affect multiple function values. sample-and-aggregate partitions the dataset into smaller subdatasets, evaluates the function of interest on each subdataset, and then aggregates the function values in a differentially private manner (e.g., using smooth sensitivity applied to the aggregation function). A single person’s data will be in only one of the subdatasets and so a single person can only affect one of the values. The advantage of the sample and aggregate approach is that requires no structural knowledge about the function of interest and is computationally efficient. (Intuitively, it pushes the privacy analysis onto the aggregation function, rather than the function we want to evaluate.) The downside is that we evaluate the function on the smaller subdatasets. This can lead to significant loss in accuracy relative to (non-privately) evaluating the function on the whole dataset. Our algorithm addresses the downside of sample and aggregate by allowing us to evaluate the function on larger subdatasets, at the expense of requiring us to evaluate on more of these subdatasets.
@@ 2.3 Lower Bounds
[AJMR15] prove lower bounds on query complexity for local Lipschitz filters. Local Lipschitz filters can be used to construct differentially private algorithms, but their lower bound does not immediately translate into a lower bound for differentially private algorithms.
@@ 3.2 Differential Privacy
Since we define differential privacy with respect to addition or removal of one person’s data, rather than replacement, the true size of the dataset $|x|$ is itself sensitive. We avoid this annoying technicality by having a separate parameter $n$ (which should be thought of as the public size of the dataset, rather than its true size) that is provided to the algorithm separately from the sensitive dataset. [^5] Alternatively, we can expend some of the privacy budget to compute a public size $n$ from the sensitive dataset. If the true size of the sensitive dataset is larger than the public size $n$, we can discard elements. If the true size of the sensitive dataset is larger than the public size $n$, we pad with nulls ($\bot$).